A scuba diver creates a spherical bubble with a radius of 3.5 cm at a depth of 30.0 m where the total pressure (including atmospheric pressure) is 4.00 atm.

What is the radius of the bubble when it reaches the surface of the water? (Assume atmospheric pressure to be 1.00 atm and the temperature to be 298 K.)

Express your answer using two significant figures.
Answer #1

We know that for a perfect gas

PV = nRT

Where
n = number of moles ( which is constant here )
R = gas constant
T = 298 K
For the two cases PV = P'V'

P = Pressure at a depth = 4 atm

P' = pressure on the surface = 1 atm

V = Volume of the sphere = (4/3)πr3

Where r = radius = 3.5 cm

So V = 179.67 cm3

V' = Volume of the sphere at the surface =?

Plug the values we get V' = PV / P'

                                       = 718.68 cm3

So V' = (4/3) πr'3

(4/3) πr'^3 = 718.68 cm3

   r'3 = 171.66

     r' = 5.557 cm

        ~5.6 cm